# Structure Your Data: Towards Semantic Graph Counterfactuals

### v1

@@ Abstract
Counterfactual explanations (CEs) based on concepts are explanations that consider alternative scenarios to understand which high-level semantic features contributed to particular model predictions. In this work, we propose CEs based on the semantic graphs accompanying input data to achieve more descriptive, accurate, and human-aligned explanations. Building upon state-of-the-art (SoTA) conceptual attempts, we adopt a model-agnostic edit-based approach and introduce leveraging GNNs for efficient Graph Edit Distance (GED) computation. With a focus on the visual domain, we represent images as scene graphs and obtain their GNN embeddings to bypass solving the NP-hard graph similarity problem for all input pairs, an integral part of the CE computation process. We apply our method to benchmark and real-world datasets with varying difficulty and availability of semantic annotations. Testing on diverse classifiers, we find that our CEs outperform previous SoTA explanation models based on semantics, including both white and black-box as well as conceptual and pixel-level approaches. Their superiority is proven quantitatively and qualitatively, as validated by human subjects, highlighting the significance of leveraging semantic edges in the presence of intricate relationships. Our model-agnostic graph-based approach is widely applicable and easily extensible, producing actionable explanations across different contexts.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Examples where semantic graphs trump concept sets. Example 1 (top) illustrates the importance of the multiplicity of concepts for edit distance and example 2 (bottom) emphasizes the intricacy of relations between them. Edits (substitutions, insertions, deletions) are enclosed in striped rectangles. Unsafe driver image sourced from (Deccan Chronicle, 2016).
@@
This work serves as an advancement of the previously presented method by Dervakos et al. (2023). We not only employ graphs to structure semantic information but also leverage Graph Neural Networks (GNNs) for the efficient approximation of GED between graph instances to compute CEs. Our findings confirm the significance of correctly representing the number and interactions of concepts and our method significantly narrows the gap to the golden standard GED, achieving closer proximity with fewer edits. To underscore the efficacy of our approach against methods with access to the underlying model, we expanded the human survey from SC and compared our CEs with the white-box CE method by Vandenhende et al. (2022) (CVE), which, despite being pixel-level, emphasizes preserving semantic consistency. Our evaluation aimed to assess both human preferences and their capacity to comprehend and anticipate the classifier’s output.
@@
We prove that our work surpasses prior SoTA approaches. Our improvements have been quantitatively assessed and further substantiated by human surveys - a significant XAI evaluation tool. The compared methodologies are diverse in terms of reliance on the features classifiers exploit and the granularity of information. To combat the challenges of evaluating the explanations, we establish unified quantitative and qualitative metrics, applicable in all cases. Further elaboration is available in the Evaluation section of §4.
Our key contributions are: 1) Demonstrating quantitative and qualitative superiority over previous black- and white-box methods, paired with enhanced adaptability due to our model-agnostic design, 2) Offering more interpretable, expressive, and actionable CEs using semantic graphs, 3) Achieving efficient GED approximation via GNNs without compromising the representation of concept interactions. Our method is novel in employing graphs and GNNs for counterfactual retrieval. We validate our approach across four diverse datasets (images & audio), using three neural and one non-neural classifier, including two human surveys and four quantitative and qualitative experiments, yielding superior outcomes and time efficiency relative to SoTA.
@@ 2 Related Work
of visual classifiers encompass pixel-level edit methods which focus on marking and altering significant image areas that influence the model’s predictions (Goyal et al., 2019; Vandenhende et al., 2022; Augustin et al., 2022). Contrary to other feature extraction counterfactual methods, the Counterfactual Visual Explanations (CVE) of Vandenhende et al. (2022) attempt to enforce semantically consistent area exchanges through an auxiliary semantic similarity component between local regions. Their semantically driven approach is the prime choice for pixel-level comparison, also providing a benchmark in contrast to techniques with direct classifier access, a feature distinguishing it from ours. Another vein of research focuses on human-interpretable concept edits to retrieve CEs. Abid et al. (2022) propose conceptual CEs in the event of a misclassification, using a white-box technique based on Concept Activation Vectors. The work by Dervakos et al. (2023) (SC) serves as an extension of Filandrianos et al. (2022), additionally emphasizing the importance of the Explanation Dataset, and expanding upon the original graph bipartite matching CE framework by leveraging the roles between concepts. Our work directly enhances these approaches, retaining advantageous qualities such as their model-agnostic nature and definition of object/relation distance through ontologies. Instead of leveraging Set Edit Distance and ignoring edges altogether (Filandrianos et al., 2022) or rolling up the edges into concepts, thus sacrificing crucial object relation information (Dervakos et al., 2023), we use the more accurate GED. To this end, we introduce semantic graphs for increased expressivity and use GNNs to accelerate GED calculation. Much like its predecessors, this paper is centered on the visual domain and applied to other modalities as a use case. Consequently, we do not delve into the literature of audio CEs, for instance. Since we propose utilizing semantic graphs corresponding to non-graph input data for CE computation, comparison to GNN explainers (Bajaj et al., 2021; Lucic et al., 2022) is not applicable. It is noteworthy that no existing method leverages GNNs for post-hoc CEs, a novel feature of our approach.
@@ 4 Experiments
comprises quantitative metrics, as well as human-in-the-loop experiments. Quantitative results are extracted by comparing the ranks retrieved based on our obtained graph embeddings to the ground truth ranks retrieved by GED. This type of analysis is not present for SC, despite its significance for objectively assessing CEs beyond intuitive metrics. The reported metrics are: 1) average Precision@k (P@k): all top-k GED retrieved results are considered relevant, 2) binary P@k and binary NDCG@k: only top-1 GED result is relevant and its position in retrieved ranks is emphasized through NDCG, 3) average number of edits: average number of node/edge insertions, deletions, and substitutions with different concepts, calculated post-hoc through GED to ensure fairness. The use of GED rankings as the golden standard for evaluation is clearly motivated by previous work (SC) and reinforced by features like: a) its purely semantic nature, b) completeness in distance representation due to its reliance on graphs which accurately encompass both objects and relations, c) deterministic nature and applicability regardless of modality and granularity of the technique under evaluation. Wide applicability is especially important because baselines include pixel-based methods which define units of information differently (significant rectangular areas vs concepts).
Human evaluation highlights several aspects of our contributions. First, to validate the quality of our retrieved CEs against SoTA, we ask our evaluators (student volunteers of engineering backgrounds) to select among two CE alternatives of a query image; an image retrieved from our method versus an image retrieved either by SC or CVE. We also test the understandability of our CEs by replicating the machine-teaching human experiment of CVE, adjusted to accommodate our graph-based explanations. We design the same stages (pre-learning, learning, and testing) and equally divide our annotators into two independent learning stage variants, namely ’visually-informed’ and ’blind’. The ’blind’ variant is the only different setting from CVE: the annotators of the ’blind’ learning stage are only provided with scene graph pairs and graph edits but no images. This evaluation method is being used for the first time to measure the reliance of humans on graph concepts rather than visual cues to understand the reasoning for classification. More information about human evaluation is provided in §A.
@@
Our presented results involve $p\sim N/2$ training graph pairs and the GCN variant unless mentioned otherwise. We produce graph representations using a single Tesla K80 GPU, while all other computations are done on CPU. We utilize PyG (Fey & Lenssen, 2019) for the implementation of GNNs and DGL (Wang et al., 2019) for approximate GED label calculation. Comparison with CVE showcases the abilities of our model-agnostic method compared to theirs, which requires white-box model access and relies on pixel-level edits. On the other hand, comparison with SC demonstrates the power of graph representations compared to set-level edits in the black-box conceptual setting. An important clarification is that SC proposes the use of roles only in the corresponding experiments of §4.3, meaning that for §4.1, 4.2 they solely rely on concepts. More details in Appendices §C, D, E.
@@ 4.1 Counterfactuals on CUB
We examine the agreement between the counterfactuals $I^{\prime}_{(B)}$ retrieved by each method (CVE, SC, ours) and the ground truth GED. Our approach outperforms CVE on every ranking metric (Tab. 2). As for SC, metrics are only valid for $k=1$ since it produces a single CE instead of a rank. Therefore, P@1 for SC is $0.022$, much lower than ours. In addition, we observe
Table 1: Average number of node, edge & total edits on CUB. Bold for best results.
[…]
@@
that our approach leads to the lowest number of overall edits: In Tab. 1, we can see that our method produces about 1 and 2 fewer edits on average for SC and CVE respectively, strengthening the claim that our CEs correspond to minimum-cost edits.
P@$\uparrow$ P@k (binary)$\uparrow$ NCDG@k (bin.$\uparrow$
k=1 k=4 k=1 k=4 k=1 k=4
CVE 0.019 0.103 0.019 0.112 0.110 0.264
Ours 0.194 0.341 0.194 0.488 0.227 0.361
@@
Figure 3: Results for Rusty $\rightarrow$ Brewer Blackbird. Color denotes best results (lower values).
@@
Analyzing the results from the comparative human survey (Tab. 4), we deduce that our CEs are more human-interpretable than both SC and CVE by a landslide: annotators prefer ours at nearly twice the rate of the CVE alternative. Compared to SC, despite the increased amount of undecided annotators, our CEs were preferred 2.6 times more frequently. This proves that despite the closeness of the two conceptual methods, ours is more intuitive to humans, confirming the meaningful addition of linking concepts within a graph. As for the machine teaching experiment, we obtain the test set accuracy scores (Tab. 4), as the ratio of correctly human-classified test images over the total number of test images. Our visually-informed accuracy clearly outperforms reported CVE scores, highlighting that concept-based CEs
Table 3: Human preference; Win%=times our method was preferred, Losefor vice-versa, Tiewhen equally preferred.
Table 4: Human test accuracy scores for correct classification of samples in classes A and B.
@@
Experiment Test acc.$\uparrow$
@@
Table 4: Human test accuracy scores for correct classification of samples in classes A and B.
are more powerful in guiding humans towards understanding discriminative concepts between classes compared to non-conceptual pixel-level CEs. The ”blind” results show an expected decrease compared to the visually-informed one, but still outperform CVE ones. The higher accuracy of concept-based over visual CEs affirms the significance humans place on higher-level features for classification.
@@
CVE may lead to non-actionable CEs, despite training on visual semantic preservation. To elaborate, we observe the following: CVE suggests that only adding a striped pattern in a Gray Catbird’s wing is adequate to classify it as a Mockingbird. However, by exhaustively generating all annotated attribute combinations of this new bird instance, we easily find several occurring attribute pairs that are not representative of the Mockingbird class; namely, no other Mockingbird has an eyering head pattern and grey breast color. Actionability dictates the prescription of attainable goals achieved through CEs that accurately represent the underlying data distribution (Poyiadzi et al., 2020). To this end, our approach not only selects CEs drawn from the existing target class distribution but also considers all edits needed to convert query to counterfactual image. Therefore, through GED we formalize a more holistic approach to distance and path between counterfactual pairs and simultaneously leverage relations between depicted objects, both visual (relations on the image) and semantic (relations mapped to WordNet synsets). Further analysis in §E.3.
@@ 4.2 Towards conceptual counterfactuals
We focus our analysis on conceptual counterfactuals since the previous sections exhibited the indisputable merits of such approaches against the SoTA pixel-level method of CVE. In the interest of experimenting on a less controlled dataset, we employ Visual Genome (VG) (Krishna et al., 2017), a dataset containing over 108k human-annotated scene graphs, describing scenes of multiple objects and their in-between interactions. We construct two manageable subsets of 500 scene graphs each, corresponding to $\sim$125k possible training graph pairs for our GNNs. The first subset denoted as VG-RANDOM is randomly selected, while the second one, named VG-DENSE, is chosen to favor higher graph densities and less isolated nodes to highlight the importance of object interconnections. Details are provided in §B. VG instances lack ground truth classification labels, allowing us to test our counterfactual retrieval method without the definition of a specific target class. We classify instances using a pre-trained Places365 classifier (Zhou et al., 2017), and regard as counterfactual classes the closest ones in rank.
@@
We first compare the average number of edits for our method and SC (Tab. 5). Initially, numerical results between the two methods seem similar, but upon closer inspection in conjunction with average GED results of Tab. 6, our method’s superiority is evident. VG contains concepts which are much more diverse than CUB, and despite the knowledge-based contraints we enforced during GED computation, edit distance is expected to be higher between concepts. Note that this is not true for mean GED since CUB has a higher number of average edits. To this end, our method leads to lower GED in all cases, even when number of edits is higher (VG-RANDOM).
Table 5: Average number of node, edge & total edits on VG.
@@
Regarding CE approximation to ground truth GED, results for our approach are presented in Tab. 7 denoted as GCN-70K. As for SC, P@1 is 0.246 on VG-DENSE and 0.204 on VG-RANDOM, compared to 0.248 and 0.214 retrieved by our method. GED approximation is satisfactory but close between methods due to a general agreement in CE retrieval. Despite this fact,
Table 6: Average top-1 GED (VG) for CEs when methods disagree.
VG-DENSE VG-RANDOM
@@
our approach still leads in all reported metrics, especially for VG-DENSE, displaying superiority in cases of disagreement. These findings place great significance in examining qualitative results.
Figure 4: Qualitative results: VG-DENSE (left 3 columns) and VG-RANDOM (right 3 columns).
Qualitative results
By examining counterfactual images retrieved for VG-DENSE in Fig. 4 (left), there is a clear indication that by considering the complex relations between concepts, our method achieves more detail-oriented results: in the 1st column, our approach not only retrieves an image with ’man’, ’board’, ’water’ concepts, but also the relation ’man on board’. In the 3rd column, we consider the relation of toppings and retrieve the pizza, while SC simply retrieves an image with similar concepts, (’bun’ and ’bread’ or ’meat’ and ’sausage’). Results on VG-RANDOM (Fig. 4 (right)) follow the same logic. In columns 4-5, our method retrieves the focal points of the images since it regards relations between trees and other objects. Taking into account the sparsity of the underlying graphs, however, in some cases the importance of concepts trumps the underlying structure, as in the 6th column. This fact is reflected in the elevated number of edits of our method for VG-RANDOM, yet it is not true for GED, showcasing once again the importance of semantic context. More details in §E.2.
Why GCN?
Ranking metrics of GNN models are provided in Tab. 7. Three GNN variants (GAT, GIN, GCN) are trained using $p=N/2=$70k scene graph pairs. The GCN-based variant consistently approaches GED the closest, with a binary P@4 of 49.0% and P@1 of 24.8% for VG-DENSE and slightly worse results on VG-RANDOM. This ablation study affirms using GCN for the GNN-based similarity component of our approach. GNNs can also outperform other prominent deterministic methods, like graph kernels (Grauman & Darrell, 2007). The reported findings grant us the security that our counterfactual explanations are trustworthy, even when applied to complex scene graphs.
Table 7: Ranking results on the two VG variants for various GNNs.
@@
Kernel 0.000 0.011 0.000 0.014 0.098 0.254
GIN-70K 0.030 0.065 0.030 0.068 0.222 0.357
GAT-70K 0.184 0.288 0.184 0.380 0.112 0.266
GCN-70K 0.214 0.300 0.214 0.418 0.250 0.380

### v2

@@ Abstract
Counterfactual explanations (CEs) based on concepts are explanations that consider alternative scenarios to understand which high-level semantic features contributed to particular model predictions. In this work, we propose CEs based on the semantic graphs accompanying input data to achieve more descriptive, accurate, and human-aligned explanations. Building upon state-of-the-art (SotA) conceptual attempts, we adopt a model-agnostic edit-based approach and introduce leveraging GNNs for efficient Graph Edit Distance (GED) computation. With a focus on the visual domain, we represent images as scene graphs and obtain their GNN embeddings to bypass solving the NP-hard graph similarity problem for all input pairs, an integral part of CE computation process. We apply our method to benchmark and real-world datasets with varying difficulty and availability of semantic annotations. Testing on diverse classifiers, we find that our CEs outperform previous SotA explanation models based on semantics, including both white and black-box as well as conceptual and pixel-level approaches. Their superiority is proven quantitatively and qualitatively, as validated by human subjects, highlighting the significance of leveraging semantic edges in the presence of intricate relationships. Our model-agnostic graph-based approach is widely applicable and easily extensible, producing actionable explanations across different contexts. The code is available at https://github.com/aggeliki-dimitriou/SGCE.
Counterfactual Explanations, Graph Neural Networks, Scene Graphs, Explainable Artificial Intelligence
@@ 1 Introduction
Figure 1: Examples where semantic graphs trump concept sets. Example 1 (top) shows the importance of the multiplicity of concepts for edit distance and example 2 (bottom) emphasizes the intricacy of relations. Edits (substitutions, insertions, deletions) are enclosed in striped rectangles. Images sourced from Visual Genome (Krishna et al., 2017), except unsafe driver (Deccan Chronicle, 2016).
@@
This work serves as an advancement of the previously presented method by Dervakos et al. (2023), which emphasizes leveraging semantically rich concepts to obtain CEs within model-agnostic settings, as long as the participating data instances are chosen wisely. However, their data representation lacks the proper incorporation of relationships between concepts, calling for a more intricate approach. We not only employ graphs to structure semantic information as a direct refinement of prior work, but also leverage Graph Neural Networks (GNNs) for the efficient approximation of GED between graph instances to compute CEs. Our findings confirm the significance of correctly representing the number and interactions of concepts and our method significantly narrows the gap to the golden standard GED, achieving closer proximity with fewer edits. To underscore the efficacy of our approach against methods with access to the underlying model, we expanded the human survey from SC and compared our CEs with the white-box CE method by Vandenhende et al. (2022) (CVE), which, despite being pixel-level, emphasizes preserving semantic consistency. Our evaluation aimed to assess both human preferences and their capacity to comprehend and anticipate the classifier’s output.
@@
We prove that our work surpasses prior SotA approaches. Our improvements have been quantitatively assessed and further substantiated by human surveys - a significant XAI evaluation tool. The compared methodologies are diverse in terms of reliance on the features classifiers exploit and the granularity of information. To combat the challenges of evaluating the explanations, we establish unified quantitative and qualitative metrics, applicable in all cases. Further elaboration is available in the Evaluation section of §4. Our key contributions are:
  • Demonstrating quantitative and qualitative superiority over previous black- and white-box methods, paired with enhanced adaptability due to our model-agnostic design,
  • Offering more interpretable, expressive, and actionable CEs using semantic graphs,
  • Achieving efficient GED approximation via GNNs without compromising the representation of concept interactions.
Our method is novel in employing graphs and GNNs for counterfactual retrieval. We validate our approach across four diverse datasets (images & audio), using three neural and one non-neural classifier, including two human surveys and four quantitative and qualitative experiments, yielding superior outcomes and time efficiency relative to SotA.
@@ 2 Related Work
of visual classifiers encompass pixel-level edit methods which focus on marking and altering significant image areas that influence the model’s predictions (Goyal et al., 2019; Vandenhende et al., 2022; Augustin et al., 2022). Contrary to other feature extraction counterfactual methods, the Counterfactual Visual Explanations (CVE) of Vandenhende et al. (2022) attempt to enforce semantically consistent area exchanges through an auxiliary semantic similarity component between local regions. Their semantically driven approach is the prime choice for pixel-level comparison, also providing a benchmark in contrast to techniques with direct classifier access, a feature distinguishing it from ours. Another vein of research focuses on human-interpretable concept edits to retrieve CEs. Abid et al. (2022) propose conceptual CEs in the event of a misclassification, using a white-box technique based on Concept Activation Vectors. The work by Dervakos et al. (2023) (SC) serves as an extension of Filandrianos et al. (2022), additionally emphasizing the importance of the Explanation Dataset, and expanding upon the original graph bipartite matching CE framework by leveraging the roles between concepts. Our work directly enhances these approaches, retaining advantageous qualities such as their model-agnostic nature and definition of object/relation distance through ontologies. Instead of leveraging Set Edit Distance and ignoring edges altogether (Filandrianos et al., 2022) or rolling up the edges into concepts, thus sacrificing crucial object relation information (Dervakos et al., 2023), we use the more accurate GED. To this end, we introduce semantic graphs for increased expressivity and use GNNs to accelerate GED calculation. A preliminary GNN-based counterfactual analysis was thoroughly discussed in the concurrent work of (Dimitriou et al., 2024), which compares various Graph Machine Learning algorithms. Much like its predecessors, our current paper is centered on the visual domain and applied to other modalities as a use case. Consequently, we do not delve into the literature of audio CEs, for instance. Despite the vast literature on graph CEs (Prado-Romero et al., 2024), comparison to GNN explainers (Bajaj et al., 2021; Lucic et al., 2022) is not applicable here, since we propose utilizing semantic graphs corresponding to non-graph input data for CE computation. It is noteworthy that no existing method leverages GNNs for post-hoc CEs, a novel feature of our approach.
@@ 4 Experiments
comprises quantitative metrics, as well as human-in-the-loop experiments. Quantitative results are extracted by comparing the ranks retrieved based on our obtained graph embeddings to the ground truth ranks retrieved by GED. This type of analysis is not present for SC, despite its significance for objectively assessing CEs beyond intuitive metrics. The reported metrics are: 1) average Precision@k (P@k): all top-k GED retrieved results are considered relevant, 2) binary P@k and binary NDCG@k: only top-1 GED result is relevant and its position in retrieved ranks is emphasized through NDCG, 3) average number of edits: average number of node/edge insertions, deletions, and substitutions with different concepts, calculated post-hoc through GED to ensure fairness.
The use of GED rankings as the golden standard for evaluation is clearly motivated by previous work (SC) and reinforced by features like: a) its purely semantic nature, b) completeness in distance representation due to its reliance on graphs which accurately encompass both objects and relations, c) deterministic nature and applicability regardless of modality and granularity of the technique under evaluation. Wide applicability is especially important because baselines include pixel-based methods which define units of information differently (significant rectangular areas vs concepts). Thus, investigating the effectiveness of GED in depth or comparing it with other similar metrics would divert from the paper’s main focus, as it is already accepted in the research community.
Human evaluation highlights several aspects of our contributions. First, to validate the quality of our retrieved CEs against SotA, we ask our evaluators (student volunteers of engineering backgrounds) to select among two CE alternatives of a query image; an image retrieved from our method versus an image retrieved either by SC or CVE. We also test the understandability of our CEs by replicating the machine-teaching human experiment of CVE, adjusted to accommodate our graph-based explanations. We design the same stages (pre-learning, learning, and testing) and equally divide our annotators into two independent learning stage variants, namely ’visually-informed’ and ’blind’. The ’blind’ variant is the only different setting from CVE: the annotators of the ’blind’ learning stage are only provided with scene graph pairs and graph edits but no images. This evaluation method is being used for the first time to measure the reliance of humans on graph concepts rather than visual cues to understand the reasoning for classification. More information about human evaluation is provided in App. A.
@@
Our presented results involve $p\sim N/2$ training graph pairs and the GCN variant unless mentioned otherwise. We produce graph representations using a single Tesla K80 GPU, while all other computations are done on a 12-core Intel Core i7-5930K CPU. We utilize PyG (Fey & Lenssen, 2019) for the implementation of GNNs and DGL (Wang et al., 2019) for approximate GED label calculation. Comparison with CVE showcases the abilities of our model-agnostic method compared to theirs, which requires white-box model access and relies on pixel-level edits. On the other hand, comparison with SC demonstrates the power of graph representations compared to set-level edits in the black-box conceptual setting. An important clarification is that SC proposes the use of roles only in the corresponding experiments of §4.3, meaning that for §4.1, 4.2 they solely rely on concepts. More details in Appendices C, D, E.
@@ 4.1 Counterfactuals on CUB
We examine the agreement between the counterfactuals $I^{\prime}_{(B)}$ retrieved by each method (CVE, SC, ours) and the ground truth GED. Our approach outperforms CVE on every ranking metric (Tab. 1). As for SC, metrics are only valid for $k=1$ since it produces a single CE instead of a rank. Therefore, P@1 for SC is $0.02$, much lower than ours. In addition, we observe that our approach leads to the lowest number of overall edits: In Tab. 2, we can see that our method produces about 1 and 2 fewer edits on average for SC and CVE respectively, strengthening the claim that our CEs correspond to the minimum number of edits.
P@$\uparrow$ P@k (binary)$\uparrow$ NCDG@k (bin.$\uparrow$
k=1 k=4 k=1 k=4 k=1 k=4
CVE 0.02 0.10 0.02 0.11 0.11 0.26
Ours 0.19 0.34 0.19 0.49 0.23 0.36
Table 2: Average number of node, edge & total edits on CUB. Bold for best results (lowest number of edits).
Node $\downarrow$ Edge $\downarrow$ Total $\downarrow$
@@
Additionally, our CEs are tied to minimum-cost edits; more specifically, the resulting GED between the query and the retrieved counterfactual scene graph obtain lower GED scores in comparison to both CVE and SC. The related analysis is presented in App. D.2.
@@
Figure 3: Results for Rusty $\rightarrow$ Brewer Blackbird. Bold denotes best results (lowest number of edits and GED scores).
@@
Analyzing the results from the comparative human survey (Tab. 3), we deduce that our CEs are more human-interpretable than both SC and CVE by a landslide: annotators prefer ours at nearly twice the rate of the CVE alternative. Compared to SC, despite the increased amount of undecided annotators, our CEs were preferred 2.6 times more frequently. This proves that despite the closeness of the two conceptual methods, ours is more intuitive to humans, confirming the meaningful addition of linking concepts within a graph. A chi-square test revealed significant differences in user preferences between our method and SC (p = 0.003) as well as our method and CVE (p = 9.21e-08), indicating a notable deviation from the expected distribution and further validating the reported results.
As for the machine teaching experiment, we obtain the test set accuracy scores (Tab. 4), as the ratio of correctly human-classified test images over the total number of test images. Our visually-informed accuracy clearly outperforms reported CVE scores, highlighting that concept-based CEs are more powerful in guiding humans towards understanding discriminative concepts between classes compared to non-conceptual pixel-level CEs. The ”blind” results show an expected decrease compared to the visually-informed ones, but still outperform CVE. The higher accuracy of concept-based over visual CEs affirms the significance humans place on higher-level features for classification. Details regarding human evaluation are presented in App. A.
Table 3: Human preference; Win%=times our method was preferred, Losefor vice-versa, Tiewhen equally preferred. Bold denotes higher human preference per method.
@@
Table 4: Human test accuracy scores for correct classification of samples in classes A and B. Bold for top accuracy score.
Human experiment Test accuracy $\uparrow$
@@
CVE may lead to non-actionable CEs, despite training on visual semantic preservation. To elaborate, we observe the following: CVE suggests that only adding a striped pattern in a Gray Catbird’s wing is adequate to classify it as a Mockingbird. However, by exhaustively generating all annotated attribute combinations of this new bird instance, we easily find several occurring attribute pairs that are not representative of the Mockingbird class; namely, no other Mockingbird has an eyering head pattern and grey breast color. Actionability dictates the prescription of attainable goals achieved through CEs that accurately represent the underlying data distribution (Poyiadzi et al., 2020). To this end, our approach not only selects CEs drawn from the existing target class distribution but also considers all edits needed to convert query to counterfactual image. Therefore, through GED we formalize a more holistic approach to distance and path between counterfactual pairs and simultaneously leverage relations between depicted objects, both visual (relations on the image) and semantic (relations mapped to WordNet synsets). Further analysis in App. E.3.
Global counterfactuals
in terms of graph edits require a standardized unit to be changed, in our case referring either to graph triples in a (concept-edge-concept) format or merely to concept edits as parts of graph triples. Both approaches regard the aggregation of local edits to explain the given classifier from a higher-level perspective. In the case of CUB, global CEs highly correlate with human perception: by considering the Parakeet Auklet $\rightarrow$ Least Auklet class transition, some key characteristics of the source class (such as the (’beak’, ’shape’, ’specialized’) triplet) need to be deleted, while others (such as the (’beak’, ’shape’, ’cone’) triplet) should be added. Further details in App. D.4.
@@ 4.2 Towards conceptual counterfactuals
We focus our analysis on conceptual counterfactuals since the previous sections exhibited the indisputable merits of such approaches against the SotA pixel-level method of CVE. In the interest of experimenting on a less controlled dataset, we employ Visual Genome (VG) (Krishna et al., 2017), a dataset containing over 108k human-annotated scene graphs, describing scenes of multiple objects and their in-between interactions. We construct two manageable subsets of 500 scene graphs each, corresponding to $\sim$125k possible training graph pairs for our GNNs. The first subset denoted as VG-RANDOM is randomly selected, while the second one, named VG-DENSE, is chosen to favor higher graph densities and less isolated nodes to highlight the importance of object interconnections. Details are provided in App. B. VG instances lack ground truth classification labels, allowing us to test our counterfactual retrieval method without the definition of a certain target class. We classify instances using a pre-trained Places365 classifier (Zhou et al., 2017), and regard as counterfactual classes the closest ones in rank. Specifically, we employ a pre-trained ResNet50 (He et al., 2015), as proposed in the original Places365 paper.
@@
We first compare the average number of edits for our method and SC (Tab. 5). Initially, numerical results between the two methods seem similar, but upon closer inspection in conjunction with average GED results of Tab. 6, our method’s superiority is evident. VG contains concepts which are much more diverse than CUB, and despite the knowledge-based contraints we enforced during GED computation, edit distance is expected to be higher between concepts. Note that this is not true for mean GED since CUB has a higher number of average edits. To this end, our method leads to lower GED in all cases, even when number of edits is higher (VG-RANDOM). Some extra analysis is provided in App. D.2.
Table 5: Average number of node, edge & total edits on VG. Bold denotes best results (lowest number of edits).
@@
Table 6: Average top-1 GED (VG) for CEs when methods disagree. Bold for best (lowest) GED scores for each dataset split.
VG-DENSE $\downarrow$ VG-RANDOM $\downarrow$
@@
Regarding CE approximation to ground truth GED, results for our approach are presented in Tab. 7 denoted as GCN-70K. As for SC, P@1 is 0.25 on VG-DENSE and 0.20 on VG-RANDOM, compared to 0.25 and 0.21 retrieved by our method. GED approximation is satisfactory but close between methods due to a general agreement in CE retrieval. Despite this fact, our approach still leads in all reported metrics, especially for VG-DENSE, displaying superiority in cases of disagreement. Ranking results consistent with our analysis are also obtained using GQA (Hudson & Manning, 2019), a VG variant which focuses on question-answering, as reported in App. D.3. Reported findings place great significance in examining qualitative results.
Table 7: Ranking results on the two VG variants for various GNNs. Bold numbers indicate best ranking metrics.
@@
Kernel 0 0.01 0 0.01 0.10 0.25
GIN-70K 0.03 0.07 0.03 0.07 0.22 0.38
GAT-70K 0.18 0.29 0.18 0.38 0.11 0.27
GCN-70K 0.21 0.30 0.21 0.42 0.25 0.38
Figure 4: Qualitative results (best metrics in bold): VG-DENSE (left 3 columns) and VG-RANDOM (right 3 columns).
Qualitative results
By examining counterfactual images retrieved for VG-DENSE in Fig. 4 (left), there is a clear indication that by considering the complex relations between concepts, our method achieves more detail-oriented results: in the 1st column, our approach not only retrieves an image with ’man’, ’board’, ’water’ concepts, but also the relation ’man on board’. In the 3rd column, we consider the relation of toppings and retrieve the pizza, while SC simply retrieves an image with similar concepts, (’bun’ and ’bread’ or ’meat’ and ’sausage’). Results on VG-RANDOM (Fig. 4 (right)) follow the same logic. In columns 4-5, our method retrieves the focal points of the images since it regards relations between trees and other objects. Taking into account the sparsity of the underlying graphs, however, in some cases the importance of concepts trumps the underlying structure, as in the 6th column. This fact is reflected in the elevated number of edits of our method for VG-RANDOM, yet it is not true for GED, showcasing once again the importance of semantic context. More details in App. E.2.
Why GCN?
Ranking metrics of GNN models are provided in Tab. 7. Three GNN variants (GAT, GIN, GCN) are trained using $p=N/2=$70k scene graph pairs. The GCN-based variant consistently approaches GED the closest, with a binary P@4 of 49% and P@1 of 24.80% for VG-DENSE and slightly worse results on VG-RANDOM. GCN systematically scores higher in comparison to theoretically more competent GNN alternatives, such as GIN. We attribute this finding to the importance of local neighborhood information for our small yet semantically dense graphs. Specifically, the VG graphs considered in our experiments rarely exceed 3-4 hops, as briefly demonstrated in App. E.2. GIN does not incorporate node features during aggregation resulting in a limited notion of semantic similarity. This ablation study affirms using GCN for the GNN-based similarity component of our approach. GNNs can also outperform other prominent deterministic methods, like graph kernels (Grauman & Darrell, 2007). The reported findings grant us the security that our counterfactual explanations are trustworthy, even when applied to complex scene graphs.